# Evaluating Sample Utility for Data Selection by Mimicking Model Weights

### v2

@@ 5 Experiments
We evaluate the effectiveness of Grad-Mimic using two experimental setups across datasets of varying scales and domains. We first test in a controllable setting, adding noise into sample labels (Sec. 5.1), followed by evaluations on large-scale web-crawled datasets (Sec. 5.2). Our goals are to validate the following claims in both setups:
  • Enhanced Data Efficiency: A pretrained reference model serves as a reliable guide, enabling mimic scores to prioritize samples that improve model performance.
  • Accurate Sample Identification: Grad-Mimic can detect noisy samples, construct effective filters, and automate data selection.
  • Dataset Quality Assessment: Mimic scores and their resulting filters provide reliable estimates of dataset quality.
@@
Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, improving data efficiency and closing the performance gap to reference model.
@@
Table 2: Stage 2 Results in Simulation Experiment: We report detection results using F1-score metric. Grad-Mimic consistently achieves precise detection of mislabeled samples across datasets and proposed identification methods. Notably, its performance remains high and robust regardless of the noise level.
5.1 Simulating Mislabeled Samples
Setups.
We begin with a controlled experiment by adding various levels of label noise to six image classification datasets. They are DTD [43], Flowers102 [44], STL10 [45], OxfordIIIT Pet [46], CIFAR10, and CIFAR100 [47]. We fine-tune the ViT-B/16 model [48] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples. We normalize mimic scores with a temperature of 0.5 and use a batch size of 32. We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last-layer weights as our reference to navigate training on the noisy datasets. We detail more training configurations in Appendix B.1.
After training, we identify samples in two ways: setting one over batch size ($1/b$) as our threshold and using $k$-means and GMM to cluster. Then we aggregate filter outputs across training steps using Snorkel framework [30].
@@ 5.1 Simulating Mislabeled Samples
Expected Results.
We aim to validate that a pretrained reference model’s weights can serve as a reliable guide. Moreover, we expect that mimic scores can help identify low-value, mislabeled samples.
Baselines.
We compare Grad-Mimic’s training method (Stage 1) with other approaches that use gradient information for sample prioritization. We consider: (i) Mini-batch SGD, the standard method, updates weights by averaging gradients within the mini-batch. (ii) GraNd [22] reweights samples based on their gradient norm, prioritizing data that induce greater changes. (iii) AGRA [23] computes the cosine similarity of each sample gradient to an average gradient of another random batch, then excludes outlier samples. (iv) Grad-Match [24] adapts gradients by solving a subset selection problem to identify key samples within each batch. Furthermore, we compare Grad-Mimic with Rho-Loss [15], a reference model-based approach that prioritizes training on worth learning samples, as well as influence function-based methods. Results are presented in Appendix C. Grad-Mimic outperforms both techniques while additionally offering efficiency advantages.
Stage 1 Results.
We present Grad-Mimic’s two-stage results separately. Stage 1 results, shown in Table 1, report testing accuracy across various methods and noisy datasets using both training configurations. These results demonstrate that Grad-Mimic effectively identifies and down-weights mislabeled samples during training, enhancing denoising and narrowing the performance gap with the reference model. Notably, in the full fine-tuning setup—where we only target the last layer to mimic—Grad-Mimic successfully guides weight updates for the entire model, ultimately outperforming other methods. This highlights its computational efficiency, as it avoids requiring sample gradients from the full set of model weights. Additionally, we explore the impact of reference model quality in Appendix C, further strengthening our approach. These findings validate the idea that using the direction induced by a better-performing model can guide model training and improve data efficiency.
Stage 2 Results.
Next, we demonstrate Grad-Mimic’s capability to detect mislabeled samples. We use F1-scores to report the detection results in Table 2. Grad-Mimic accurately identifies mislabeled data across datasets and remains robust when facing different noise levels.
@@
We use the ensemble filter from each dataset created by GMM clustering to estimate training dataset quality using the proportion of retained data. The results, displayed in Figure 4, highly correspond to the presence of label noise, with a Pearson correlation of -0.903. These demonstrate mimic score’s effectiveness for dataset quality estimation.
5.2 Data Selection in Large-scale Web Datasets
Setups.
Next, we test Grad-Mimic in a more challenging setting using million-scale web-crawled datasets. We use the small-scale and medium-scale DataComp datasets [17], which contain approximately 10 million and 100 million image-caption pairs [^1] We identified that some URLs provided by DataComp dataset are now broken. This means that our results might not be comparable to previous approaches on the DataComp Leaderboard. See here for details., respectively, to train CLIP model from scratch [2]. In these datasets, noise is naturally presented in the data. We follow the training setup from DataComp [49, 17] and use publicly available pretrained CLIP model’s weights as our reference  [^2] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K. This reference model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us. It serves as a proxy for the ideal reference point for scenarios where training such powerful models is infeasible. We use the final MLP layer in the text and image encoders respectively as our target to mimic. We evaluate the performance of resulting model using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks. Other details can be found in Appendix B.2.
In Stage 2, we use top-$k$ percent identification to create the final filter, subsequently training a second CLIP model on the curated dataset to validate resulting filter’s effectiveness.
@@ 5.2 Data Selection in Large-scale Web Datasets
Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic yields higher CLIP model performance. Full table with various temperature settings is placed in Table 6 in Appendix  C.
Expected Results.
We anticipate mimic scores help large-scale training focus on high-value samples and can serve as instrumental metric for automating effective data curation.
Baselines.
In the first stage, we compare Grad-Mimic to vanilla training, where each sample contributes in the weight update equally. After training, we compare our mimic score-based filter against the following methods: (1) No Filtering: using the entire training pool, (2) Basic Filtering [17]: selecting samples based on predefined criteria like caption length, image size, and caption language—representing a human-designed filter, (3) CLIP Score [18]: selecting top-$k$ percent samples based on embedding alignment between images and captions. Scores are computed by OpenAI’s CLIP ViT-B/32 model pretrained on over 400 million samples [2]. We see CLIP score as a strong baseline, as it is prebuilt through heavy training using hundreds of millions of samples.
Stage 1 Results.
Table 3 presents pretraining results for CLIP models on both dataset scales. Grad-Mimic consistently outperforms standard training across all temperature settings. Full results using different temperatures are provided in Appendix C. Interestingly, we find that mimicking the last layer of the image encoder yields more performance gains compared to targeting the text encoder. These results highlight Grad-Mimic’s success in using online pretrained weights as a reference to prioritize high-value samples for learning.
@@
((a))
((b))
Figure 4: Mimic scores offer precise dataset quality estimation: Left figure computes the portion of retained samples to evaluate dataset quality in simulation experiments. Right figure takes the average on mimic scores in each filtered dataset for estimation. [S/M] denotes the scale of DataComp dataset—small [S] or medium [M]—that the filter is designed on.
Stage 2 Results.
We use the derived mimic scores from our best-performing model in Stage 1 to design filters and evaluate their effectiveness. Results are presented in Table 4. Grad-Mimic automates the selection, outperforming basic filtering strategy at both scales. Additionally, we remove low-value samples identified by mimic score to improve CLIP score-based filter (B/32 Top 30%). This complementary approach not only enhances model performance but also improves data efficiency by reducing the training set size, specifically cutting down 4.7 million samples in the medium-scale dataset.
Lastly, we gather mimic scores from datasets that are curated by various CLIP score-based filters. We include the top-ranked approach during our time of experimentation [50], which uses an orthogonal approach based on object detection models. We average samples’ mimic scores in each curated dataset and compare with their corresponding performance gains. The results, shown in Figure 4, reveal a positive alignment with a Pearson correlation of 0.958, demonstrating mimic score as a reliable metric for dataset quality estimation.
Sample Analysis.
We analyze high-value and low-value samples ranked by mimic score. In Figure 5, we compare token usage in captions. We find that bottom-ranked samples often carry captions that are either too short (1-2 words) or excessively long (over the maximum token limit), unlike top-ranked samples. Moreover, mimic score-based filter effectively identifies low-value samples, such as Captcha-like images in the small-scale dataset and empty or incoherent images in the medium-scale dataset (see examples in Figure 1). These filtering patterns align with human intuition but are automatically captured by Grad-Mimic.
We visualize the top and bottom 30% of samples along with their CLIP scores in Figure 6. High-value samples align with higher CLIP scores, while low-value ones are more concentrated in the lower range (half below 0.2). Furthermore, we calculate their correlation on random samples. We normalize both scores into the range 0 to 1 for clarity The observed positive correlation supports mimic scores in identifying high-quality samples.
Figure 5: The filtering logic behind Grad-Mimic aligns with handcrafted heuristics, particularly in caption length. Low-value samples (left figure) tend to have captions that are either too short or exceed the maximum token length, i.e., 77.
((a))
((b))
Figure 6: Mimic score distribution and correlation to CLIP score.
@@
DataComp’23 Top-ranked 0.166 0.284 0.271 0.426
Predicting Pretraining Dataset
We explore an interesting application of mimic score: predicting the pretraining dataset. Specifically, we ask: how accurately can Grad-Mimic check whether a given sample was used to train a reference model? We test this hypothesis on the small-scale DataComp dataset. We first apply various pre-built filters to curate datasets and train CLIP models on each. Then, we target their final-layer weights to mimic and apply Grad-Mimic to train on the entire data pool. Our goal is to see whether the top-ranked samples (matched in size to the curated datasets) identified by mimic score appear in their training dataset.
We evaluate our approach against random selection using Jaccard similarity and percentage of overlap. As shown in Table 5, mimic score-based selection identifies more samples used to train the reference model compared to random sampling. We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset by simply mimicking final-layer weights without direct access to their filtering steps.
@@ 6 Conclusion
We introduce Mimic Score, an efficient data quality metric that assesses samples’ utility in model training. We quantify samples by measuring the alignment between gradients and the vector toward pretrained reference weights. Using this, we develop Grad-Mimic to automate data selection. Empirically, Grad-Mimic effectively prioritizes samples for learning, improving data efficiency and model performance. Moreover, Mimic Score aids in creating effective filters and estimating dataset quality. Resulting filters outperform human-designed ones and exhibit a strong correlation with dataset quality and performance gains.
7 Impact Statements
Efficient data selection is increasingly important for training foundation models. We demonstrate Grad-Mimic’s effectiveness in identifying and removing low-value samples, which enhances data efficiency and resulting model performance. We do not foresee explicit negative social impacts arising from Grad-Mimic. However, if the reference model used to evaluate sample utility is suboptimal, there remains a risk of refining low-quality datasets (we have an analysis of the impact of reference model quality in Appendix C). We believe that careful selection and validation of this reference model, a key area of future research, will minimize this risk and enable the creation of improved datasets.

### v3

@@ 5 Experiments
We validate Grad-Mimic’s effectiveness using two experimental setups across diverse scales and domains. First, we test in a controlled setting, adding noise sample labels (Sec. 5.1), then evaluate on large-scale web-crawled datasets (Sec. 5.2). Our goals are to confirm key claims in both settings:
  • Enhanced Data Efficiency: The weights of pretrained reference model serve as reliable guide. Using derived mimic scores to prioritize samples improves data efficiency and speed up convergence.
  • Accurate Sample Identification: Grad-Mimic precisely detects noisy samples, constructs effective ensemble filter, and automates data selection for high-quality datasets.
  • Dataset Quality Assessment: Mimic scores and their resulting filters provide reliable estimates of dataset quality and predict model performance gains.
5.1 Simulating Mislabeled Samples
Setups. We begin with a controlled experiment by adding various levels of label noise to six image classification datasets. They are DTD [48], Flowers102 [49], STL10 [50], OxfordIIIT Pet [51], CIFAR10, and CIFAR100 [52]. We fine-tune the ViT-B/16 model [53] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples. We normalize mimic scores with a temperature of 0.5 and use a batch size of 32. We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last-layer weights as our reference to navigate training on the noisy datasets. We detail more training configurations in Appendix C.
Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, improving data efficiency and denoising capability.
@@ 5.1 Simulating Mislabeled Samples
Table 2: Stage 2 Results in Simulation Experiment: We report detection results using F1-score. Grad-Mimic reliably detects mislabeled samples across datasets and binarization methods, maintaining strong performance despite varying noise levels.
@@
After training, we binarize samples’ mimic scores in two ways: setting one over batch size ($1/b$) as our threshold and using $k$-means and GMM to cluster. Then we aggregate filter outputs across training steps using Snorkel framework [35].
@@
Expected Results. We aim to validate that a pretrained reference model’s weights can serve as a reliable selection guide, and we expect mimic scores can help identify low-value, mislabeled samples.
Baselines. We compare Grad-Mimic’s Stage 1 online method with other approaches that use gradient information for sample prioritization. We consider: (i) Mini-batch SGD, the standard method, updates weights by averaging gradients within the mini-batch. (ii) GraNd [27] reweights samples based on their gradient norm, prioritizing data that induce greater changes. (iii) AGRA [28] computes the cosine similarity of each sample gradient to an average gradient of another random batch, then excludes outlier samples. (iv) Grad-Match [29] adapts gradients by solving a subset selection problem to identify key samples within each batch.
Furthermore, we compare Grad-Mimic with Rho-Loss [14], a reference model-based approach that prioritizes training on worth learning samples, as well as approximated influence function-based methods. Results are presented in Appendix D. Grad-Mimic outperforms both techniques while additionally offering efficiency advantages (See Appendix E).
Stage 1 Results. We present Grad-Mimic’s two-stage results separately. Stage 1 results, shown in Table 1, report testing accuracy across various methods and noisy datasets using both training configurations. These results demonstrate that Grad-Mimic effectively identifies and down-weights mislabeled samples during training, enhancing denoising capability. Notably, in the full fine-tuning setup—where we only target the last layer to mimic—Grad-Mimic successfully guides weight updates for the entire model, ultimately outperforming other baseline methods. Beyond denoising improvements, Grad-Mimic enjoys many efficiency benefits than others, including faster convergence, lower compute runtime, and reduced GPU memory usage. We place these analysis in Appendix E. Additionally, we explore the impact of reference model quality in Appendix F, further strengthening our approach. These findings all validate the idea that using the direction induced by a better-performing model can guide model training and improve data efficiency.
Stage 2 Results. Next, we demonstrate Grad-Mimic’s capability to detect mislabeled samples. We use F1-scores to report the detection results in Table 2. Grad-Mimic accurately identifies mislabeled data across datasets and remains robust when facing different noise levels.
@@
We use the ensemble filter from each dataset created by GMM clustering to estimate training dataset quality using the proportion of retained data. The results, displayed in Figure 4 (middle), highly correspond to the presence of label noise, with a Pearson correlation of -0.903. These demonstrate mimic score’s effectiveness for dataset quality estimation.
Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic yields higher CLIP model performance. Full table with various temperature settings is placed in Table 8 in Appendix  F.
@@
5.2 Data Selection in Large-scale Web Datasets
Setups. Next, we test Grad-Mimic in a more challenging setting using million-scale web-crawled datasets. We use the small-scale and medium-scale DataComp datasets [21], which contain approximately 10 million and 100 million image-caption pairs [^2] We identified that some URLs provided by DataComp dataset are now broken. This means that our results might not be comparable to previous approaches on the DataComp Leaderboard. See here for details., respectively, to train CLIP model from scratch [3]. In these datasets, noise is naturally presented in the data. We follow the training setup from DataComp [54, 21] and use publicly available pretrained CLIP model’s weights as our reference  [^3] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K. This reference model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us. It serves as a proxy for the ideal reference point for scenarios where training such powerful models is infeasible. We use the final MLP layer in the text and image encoders respectively as our target to mimic. We evaluate the performance of resulting model using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks. Other implementation details can be found in Appendix C.
In Stage 2, we use top-$k$ percent approach to binarize scores and create the final filter, subsequently training a second CLIP model on the curated dataset to validate resulting filter’s effectiveness.
Expected Results. We anticipate mimic scores help large-scale training focus on high-value samples and can serve as instrumental metric for automating effective data curation.
Baselines. For the first stage, we compare Grad-Mimic to vanilla training, where each sample contributes in the weight update equally. For Stage 2 comparison, we compare our mimic score-based filter against the following methods: (1) No Filtering: using the entire training pool, (2) Basic Filtering [21]: selecting samples based on predefined criteria like caption length, image size, and caption language—representing a human-designed filter, (3) CLIP Score [22]: selecting top-$k$ percent samples based on embedding alignments between images and captions. Scores are computed by OpenAI’s CLIP ViT-B/32 model pretrained on over 400 million samples [3]. We see CLIP score as a strong baseline, as it is prebuilt through heavy training using hundreds of millions of samples.
Stage 1 Results. Table 3 presents pretraining results for CLIP models on both dataset scales. Grad-Mimic consistently outperforms standard training at all temperature settings. Full results using various temperatures are provided in Appendix F. Interestingly, we find that mimicking the last layer of the image encoder yields more performance gains compared to targeting the text encoder. Figure 4 (left) presents the learning curves for CLIP models trained with Grad-Mimic versus the vanilla method. By re-weighting samples based on mimic scores in weight updates, Grad-Mimic achieves faster convergence, requiring 20.7% fewer steps. These findings highlight Grad-Mimic’s success in leveraging online pretrained weights to prioritize high-value samples for efficient training.
@@ 5.2 Data Selection in Large-scale Web Datasets
Figure 4: The left figure shows mimic score-guided training converges faster than vanilla baseline. The middle one uses retention rate to estimate training dataset quality in simulation experiments. The right figure takes the average on mimic scores in each curated dataset for performance gain estimation. [S/M] denotes the scale of DataComp dataset, small [S] or medium [M], that the filter is designed on.
Stage 2 Results. We use the derived mimic scores from our best-performing model in Stage 1 to design filters and evaluate their effectiveness. Results are presented in Table 4. Grad-Mimic automates the selection, outperforming basic filtering strategy at both scales. Additionally, we remove low-value samples ranked by mimic score to improve CLIP score-based filter (B/32 Top 30%). This complementary approach not only enhances model performance but also improves data efficiency by reducing training set size, specifically removing 4.7 million samples in the medium-scale dataset.
Lastly, we gather mimic scores from datasets that are curated by various CLIP score-based filters. We include the top-ranked approach during our time of experimentation [55], which uses an orthogonal approach based on object detection models. We average samples’ mimic scores in each curated dataset to estimate dataset quality and compare with their corresponding performance gains. The results, shown in Figure 4 (right), reveal a positive alignment with a Pearson correlation of 0.958, demonstrating the feasibility to use mimic score as a reliable metric to predict potential model performance gains.
Sample Analysis. We analyze high- and low-value samples ranked by mimic score. In Figure 5, we compare token usage in captions. We find that bottom-ranked samples often carry captions that are either too short (1-2 words) or excessively long (over the maximum token limit), unlike top-ranked samples. Moreover, mimic score-based filter effectively identifies low-value samples, such as misaligned-caption images in the medium-scale dataset (see examples in Figure 1). These filtering patterns align with human intuition but are automatically captured by Grad-Mimic.
We visualize the top and bottom 30% of samples along with their CLIP scores in Figure 5. High-value samples align with higher CLIP scores, while low-value ones are more concentrated in the lower range (half below 0.2). Furthermore, we calculate their correlation on random samples. We normalize both scores into the range 0 to 1 for clarity. The observed positive correlation supports mimic scores in identifying high-quality samples (those have high CLIP score).
Figure 5: The left two figures compare token counts of bottom- and top-ranked samples by mimic score. The filtering logic matches handcrafted heuristics, particularly in caption length, where low-value samples have captions that are either too short or over the 77-token limit. The right two figures show mimic score distribution and its correlation with CLIP score.
@@
DataComp’23 Top-ranked [55] 0.166 0.284 0.271 0.426
Predicting Pretraining Dataset. We explore an interesting application of mimic score: predicting the pretraining dataset. Specifically, we ask: how accurately can Grad-Mimic check whether a given sample was used to train a reference model? We test this hypothesis on the small-scale DataComp dataset. We first apply various pre-built filters to curate datasets and train CLIP models on each. Then, we target their final-layer weights to mimic and apply Grad-Mimic to train on the entire data pool. Our goal is to see whether the top-ranked samples (matched in size to the curated datasets) identified by mimic score appear in their training dataset.
We evaluate our approach against random selection using Jaccard similarity and percentage of overlap. As shown in Table 5, mimic score-based selection identifies more samples used to train the reference model compared to random sampling. We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset [55] by simply mimicking final-layer weights without direct access to their filtering steps.
@@ 6 Conclusion
We introduce Mimic Score, an efficient data quality metric that assesses samples’ utility in model weight updates. By measuring the alignment of gradients with the vector toward pretrained reference weights, Mimic Score enables precise data assessment. Leveraging this scoring metric, we develop Grad-Mimic, a two-stage framework that automates high-quality data selection. Theoretically, we study how mimic score-guided training can converge faster than gradient descent. Empirically, Grad-Mimic effectively improves data efficiency, accelerating model convergence, and accurately detect mislabeled samples. The resulting filters outperform human-designed ones, and their statistics exhibit a strong correlation to training dataset quality and model performance improvements.